alones (alones):

Hm._. I keep on getting the wrong one /:

alones (alones):

Evaluate m0 - n2 for m = 2 and n = -1. But I get 2

OpenStudy (onepieceftw):

Is the m0 - n2 actually m^0 = n^2?

OpenStudy (shadowlegendx):

hi

OpenStudy (onepieceftw):

m^0 - n^2*

alones (alones):

yes FTW SHAD I WAS ABOUT TO TAG YOU LOL

OpenStudy (onepieceftw):

Okay. well for 2^0 it equals 1. Because anything to the power of zero equals 1.

OpenStudy (onepieceftw):

and -1^2 = 1

alones (alones):

Oh okay thank you ;)

OpenStudy (onepieceftw):

You're welcome.